OpenStudy (anonymous):

The combined area of two squares is 250 square feet. Each side of one square is three times as long as a side of the other square. What would be the lengths of the sides of each square.

OpenStudy (lgbasallote):

A_1 + A_2 = 250 A_1 = x^2 A_2 = (3x)^2 x^2 + (3x)^2 = 250 i leave the rest to you :)

OpenStudy (campbell_st):

draw a picture  area of the small square is A = x^2 Area of the large square is A = (3x)^2 A = 9x^2 the sum of the areas is 250 250 = x^2 + 9x^2 250 = 10x^2 25 = x^2 x = 5 small square has side length 5 ft and large 15ft